# A Precision Gyroscope from the Spin of Light

### v1

@@ Abstract
We describe a gyroscope that measures rotation based on the effects of the rotation on the polarization of light. Rotation induces a differential phase shift in the propagation of left and right circularly polarized light and this phase shift can be measured in suitably designed interferometric setups. The signal in this setup is independent of the frequency of light, unlike various sources of noise such as vibrations, which cause phase shifts that depend on the frequency. Such vibrations are the practical limit on the sensitivity of conventional Sagnac-style optical interferometers that are typically used as gyroscopes. In the proposed setup, one can potentially mitigate this source of noise by simultaneously using two (or more) sources of light that have different frequencies. The signal in this setup scales with the total storage time of the light. Due to its frequency independence, it is thus most optimal to measure the signal using superconducting RF systems where the high finesse of the available cavities enables considerably longer storage times than is possible in an optical setup.
@@ I Introduction
The basic physical principle behind this measurement scheme is as follows. Rotation changes the dispersion relation of circularly polarized light (see e.g., [8, 9, 10, 11, 12, 13, 14]), leading to a differential shift in the dispersion relation for left- and right-circularly polarized light. This shift leads to a relative phase between suitably stored right- and left-circularly polarized light, which can be measured precisely using interferometry. The key advantage of this method of measurement, as we will show below, is that the phase shift is independent of the frequency of light: it depends only on the rotation rate and the time for which the light is stored. This is unlike the case of the Sagnac effect that is typically used to measure rotation using light, where the rotation manifests as a length change. In that case, the phase shift is directly proportional to the frequency of light. A key limitation of gyroscopes based on the Sagnac effect is vibration of the mirrors used in the setup; such vibrations cause relative length changes in the apparatus which lead to phase shifts that are also directly proportional to the frequency of light, preventing the system from distinguishing between rotations and vibrations. However, in the measurement scheme proposed in this paper, the phase shift caused by polarization is independent of the frequency of the light whereas phase shifts arising from vibrations will be directly proportional to the frequency. Due to this difference, one can envisage a setup where two frequencies of light are used to measure simultaneously, breaking the degeneracy between the signal from rotation and the noise from vibration. That is, because rotations and vibrations cause phase shifts that scale differently with frequency, the phase shift difference between the two frequencies can be used to measure and subtract out vibrations, giving access to the signal from rotations that is common to both frequencies.
@@
The rest of this paper is organized as follows. In section II, we compute the effects of rotation on left- and right-circularly polarized light. Following this, in section III we propose a measurement scheme to detect this effect in a suitably designed Fabry–Pérot cavity. We estimate the fundamental sensitivity of such a scheme in section IV and compare the reach of this scheme to that of the conventional Sagnac scheme. We also comment on various systematic effects that may prevent us from achieving this fundamental limit and suggest some potential ways to mitigate these effects. Following this, we conclude in section V. Appendix A gives more general discussion of solutions for light propagating in the assumed rotating cavity setups than those discussed in the main text. The analysis in the main text of this paper is performed in a rotating frame; we describe the analysis in the inertial frame in appendix B.
@@ II The Signal
$ds^{2}=\left(-1+\Omega^{2}\left(x^{2}+y^{2}\right)\right)dt^{2}-2\,\Omega\,y\,dtdx+2\,\Omega\,x\,dtdy+dx^{2}+dy^{2}+dz^{2}\>.$ (1)
@@
$(A_{\pm})_{\mu}=\mathcal{A}_{\pm}^{0}e^{i\left(\omega t-k_{\pm}z\right)}\left(0,1,\pm\,i,A_{z}\left(t,x,y,z\right)\right)\>,$ (2)
where $\mathcal{A}_{\pm}$ correspond, respectively, to right (--) and left ($+$) circularly polarized light. [^1] Given our sign conventions, the left/right definitions for handedness adopted here match those seen by the emitter. Here, $A_{z}=\pm i\Omega(x\pm iy)$, which vanishes in the limit $x\rightarrow 0$, $y\rightarrow 0$ (appropriate when the transverse size of the light beam is small compared to $1/\Omega$). Waves of the form at (2) are solutions to the vacuum Maxwell equations that have the following dispersion relations:
@@
This shows that right- and left-circularly polarized light will travel with different wave-numbers $k_{\pm}$. If they travel for a distance $L$, the differential phase between the right and left circularly polarized light will be $\Omega L$, independent of the frequency $\omega$ of the light. We discuss more general orientations for the rotation with respect to the propagation direction of the light in appendix A.
@@ III The Setup
The optimal way to measure the shift (3) in the wavenumber caused by the rotation is to convert it into a phase. To maximize this phase, it is advantageous to expose the light to the rotation for as long as possible. In order to achieve this in a compact device, the light has to be held in a cavity of some kind. Further, we also need a measurement scheme that significantly suppresses dominant sources of noise such as frequency and phase noise in the light, as well as systematics that may arise from vibrations in the cavity. Motivated by these considerations, we propose the setup described in figure 1.
@@
Figure 1: The Scheme to Measure Rotation: Linearly polarized light is sent through a suitable polarization apparatus which splits the beam into left (red) and right (green) circularly polarized beams. These are stored in two Fabry–Pérot cavities that are aligned with the axis of rotation $\vec{\Omega}$. The differential phase between these two beams is free of phase and frequency noise from the light source, but retains the rotation signal.
@@
In this setup, we have two Fabry–Pérot cavities, each of length $L$ and finesse $\mathcal{F}$. The cavities are aligned along the direction of rotation $\vec{\Omega}$. We produce linearly polarized light from a suitable source. This light is split and one part of it is sent to a suitable optical element [i.e., a quarter-wave ($\lambda/4$) plate] which converts it to left circularly polarized light while the other part is sent to a different quarter-wave plate that converts that part to right circularly polarized light. The two beams are sent to two different Fabry–Pérot cavities where they are stored for a time $\sim\mathcal{F}L$. After this time, the beams are sent back through the quarter-wave plates and recombined. The interfered beam is detected at the detector.
@@
In this scheme, the wavenumbers of the left- and right-circularly polarized light experience different shifts. This difference manifests as a relative phase shift between these two beams when they are re-interfered. We want the signal in this setup to scale with $\mathcal{F}L$, the total time for which the light is held in the cavity. If the mirrors of the cavity are made with conventional reflecting surfaces, the overall phase of the electromagnetic vector potential $\left(A_{\pm}\right)_{\mu}$ changes by $\pi$ upon reflection. But, the reflection does not cause a relative phase between the $x$ and $y$ components of $\left(A_{\pm}\right)_{\mu}$. It can then be shown that for the reflected wave also the wave-numbers are $k_{\pm}=\omega\mp\Omega$. Thus, the phase accumulated by the light will continually add. In other words, while left- (right-) circularly polarized light becomes right- (left-) circularly polarized upon reflection, since the direction of the light has also changed, the effect of the rotation on the wave-number remains the same. [^2] The situation is thus similar to the accumulation of phase for light stored in a cavity in the presence of a static axion gradient (cf. [15]), as opposed to a long-period time-dependent dark-matter axion field for which the phase accumulation cancels out when employing conventional mirrors in the cavity (cf. [12]). Another way to understand the continual addition of phase regardless of the direction of travel of the light in the cavity is that the flip of left- to right-handedness (and vice versa) of the light upon reflection off the mirrors is an helicity flip that occurs because the linear momentum of the light changes sign but the angular momentum does not. Because the relative orientation of the angular momentum of the light and that of the rotation of the cavities is however unchanged after the bounce off the mirror, the phase shift to the light induced by the rotation continues to add for trips in either direction in the cavity.
@@ IV Sensitivity
The envisaged operation of this light-based gyroscope would be as a final fine-measurement stage. That is, we presume that this gyroscopic assembly would be mounted on a set of nulling gimbals designed to cancel most of the gross rotational motion of the platform being monitored. It is for this reason that we have analysed only rotational motion oriented around the axis of the light propagation; large rotational motion around another axis would induce additional phase shifts and phase gradients across cavity mirrors that would complicate the analysis. If fine measurement along multiple axes is required, either a set of orthogonal light gyroscopes could be used, or the additional phase shifts induced by rotation about an axis different from the cavity symmetry axis would need to be understood in more detail. See also appendix A for further discussion.
@@ Acknowledgements.
Appendix A Generic rotations
Our purpose in this appendix is to demonstrate that the phase shift of the transverse vector potentials considered in the main text still holds even when there are additional small rotations about axes orthogonal to the cavity symmetry axis.
@@ Appendix A Generic rotations
we can perturb about the solution at (2), which is exact when $\Omega_{x}=\Omega_{y}=0$. Suppose that we treat $\Omega_{x},\Omega_{y}$ as formal small parameters by sending $\Omega_{x,y}\rightarrow\epsilon\Omega_{x,y}$, and that we write each component of the vector potential as a power series in $\epsilon$. Substituting into Maxwell’s equations (MEs), expanding in powers of $\epsilon$, and satisfying the resulting equations that occur at each independent power of $\epsilon$ separately, allows us to derive a systematic perturbative expansion in powers of the formal small parameter $\epsilon$. Moreover, we will also consider the solution only close to the symmetry axis of the cavity, such that we can take $x,y\rightarrow 0$ after substituting into MEs. [^3] Of course, because ME are differential equations, it is important for the purposes of checking if this is a good solution to keep terms $\sim x^{n},y^{m}$ until after substitution in the MEs. That is, to only set $x,y\rightarrow 0$ after such substitution.
@@
We therefore conclude that, we long as the mild conditions derived in this appendix hold, the approach we advanced in the main text will continue to apply even when $\Omega_{x,y}\neq 0$, so long as these rotation rates are sufficiently small.
Appendix B Inertial frame and rotation of the apparatus
@@
$A_{\mu}=e^{i((\omega-\sigma\Omega)t-kz)}(0,1,i\sigma,0)\quad,\quad k=\omega-\sigma\Omega\;.$ (10)
@@
This makes clear that reading the signal from a rotating lab does not remove the polarization-dependent phase shift. Indeed, while the time dependent part of the phase-shift between different polarizations is absorbed by the basis, the space dependent relative phase accumulated over the interval $L$ does not cancel. This is the same phase-shift one sees in the non-inertial frame.
@@
$2A_{\mu}=e^{i((\omega-\Omega)t-k_{+}z)}(0,1,i,0)+e^{i((\omega+\Omega)t-k_{-}z)}(0,1,-i,0)\quad,\quad k_{\pm}=\omega\mp\Omega\;.$ (12)
This solution does again satisfy $A_{\mu}(t,z)=A_{\mu}^{B}(t-z)$, meaning that as a wavefront propagates it remains a linear combination of the two circular polarizations emitted initially. Since these have different dispersion relations, they accumulate a relative phase, resulting in a rotation of the plane of linear polarization (cf. [8, 9, 10, 11, 12, 13, 14]).

### v2

# A Precision Gyroscope from the Helicity of Light

@@ Abstract
We describe a gyroscope that measures rotation based on the effects of the rotation on the polarization of light. Rotation induces a differential phase shift in the propagation of left- and right-circularly polarized light and this phase shift can be measured in suitably designed interferometric setups. The signal in this setup is independent of the frequency of light, unlike various sources of noise such as vibrations, which cause phase shifts that depend on the frequency. Such vibrations are the practical limit on the sensitivity of conventional Sagnac-style optical interferometers that are typically used as gyroscopes. In the proposed setup, one can potentially mitigate this source of noise by simultaneously using two (or more) sources of light that have different frequencies. The signal in this setup scales with the total storage time of the light. Due to its frequency independence, it is thus most optimal to measure the signal using superconducting radio-frequency systems where the high finesse of the available cavities enables considerably longer storage times than is possible in an optical setup.
preprint: FERMILAB-PUB-24-0325-SQMS-TD
@@ I Introduction
The basic physical principle behind this measurement scheme is as follows. Rotation changes the dispersion relation of circularly polarized light (see, e.g., [8, 9, 10, 11, 12, 13, 14]), leading to a differential shift in the dispersion relation for left- and right-circularly polarized light. This shift leads to a relative phase between suitably stored right- and left-circularly polarized light, which can be measured precisely using interferometry. The key advantage of this method of measurement, as we will show below, is that the phase shift is independent of the frequency of light: it depends only on the rotation rate and the time for which the light is stored. This is unlike the case of the Sagnac effect that is typically used to measure rotation using light, where the rotation manifests as a length change. In that case, the phase shift is directly proportional to the frequency of light. A key limitation of gyroscopes based on the Sagnac effect is vibration of the mirrors used in the setup; such vibrations cause relative length changes in the apparatus which lead to phase shifts that are also directly proportional to the frequency of light, preventing the system from distinguishing between rotations and vibrations. However, in the measurement scheme proposed in this paper, the phase shift caused by polarization is independent of the frequency of the light whereas phase shifts arising from vibrations will be directly proportional to the frequency. Due to this difference, one can envisage a setup where two frequencies of light are used to measure simultaneously, breaking the degeneracy between the signal from rotation and the noise from vibration. That is, because rotations and vibrations cause phase shifts that scale differently with frequency, the phase-shift difference between the two frequencies can be used to measure and subtract out vibrations, giving access to the signal from rotations that is common to both frequencies.
@@
The rest of this paper is organized as follows. In Sec. II, we compute the effects of rotation on left- and right-circularly polarized light. Following this, in Sec. III we propose a measurement scheme to detect this effect in a suitably designed Fabry–Pérot cavity. We estimate the fundamental sensitivity of such a scheme in Sec. IV and compare the reach of this scheme to that of the conventional Sagnac scheme. We also comment on various systematic effects that may prevent us from achieving this fundamental limit and suggest some potential ways to mitigate these effects. Following this, we conclude in Sec. V. In Appendix A we give an explicit derivation of our solution for a plane wave in the rotating frame. In Appendix B we give more general discussion of solutions for light propagating in rotating cavity setups other than those discussed in the main text. The analysis in the main text of this paper is performed in a rotating frame; we describe the analysis in the inertial frame in Appendix C.
@@ II The Signal
$ds^{2}=\left[-1+\Omega^{2}\left(x^{2}+y^{2}\right)\right]dt^{2}-2\,\Omega\,y\,dtdx+2\,\Omega\,x\,dtdy+dx^{2}+dy^{2}+dz^{2}\>.$ (1)
@@
$(A_{\pm})_{\mu}=\mathcal{A}_{\pm}^{0}e^{i\left(\omega t-k_{\pm}z\right)}\big{[}0,1,\pm\,i,A_{z}\left(t,x,y,z\right)\!\big{]}\>,$ (2)
where $\mathcal{A}_{\pm}$ correspond, respectively, to right- (--) and left- ($+$) circularly polarized light. [^1] Given our sign conventions, the left/right definitions for handedness adopted here match those seen by the emitter. Here, $A_{z}=\pm i\Omega(x\pm iy)$, which vanishes in the limit $x\rightarrow 0$, $y\rightarrow 0$ (appropriate when the transverse size of the light beam is small compared to $1/\Omega$). Waves of the form at (2) are solutions to the vacuum Maxwell equations that have the following dispersion relations:
@@
We give a derivation of this result in Appendix A. This shows that right- and left-circularly polarized light will travel with different wavenumbers $k_{\pm}$. If they travel for a distance $L$, the differential phase between the right- and left-circularly polarized light will be $\Omega L$, independent of the frequency $\omega$ of the light. We discuss more general orientations for the rotation with respect to the propagation direction of the light in Appendix B.
@@ III The Setup
The optimal way to measure the shift (3) in the wavenumber caused by the rotation is to convert it into a phase. To maximize this phase, it is advantageous to expose the light to the rotation for as long as possible. In order to achieve this in a compact device, the light has to be held in a cavity of some kind. Further, we also need a measurement scheme that significantly suppresses dominant sources of noise such as frequency and phase noise in the light, as well as systematics that may arise from vibrations in the cavity. Motivated by these considerations, we propose the setup described in Fig. 1.
@@
Figure 1: The scheme to measure rotation. Linearly polarized light is sent through a suitable polarization apparatus which splits the beam into left- (red) and right- (green) circularly polarized beams. These are stored in two Fabry–Pérot cavities that are aligned with the axis of rotation $\vec{\Omega}$. The differential phase between these two beams is free of phase and frequency noise from the light source, but retains the rotation signal.
@@
In this setup, we have two Fabry–Pérot cavities, each of length $L$ and finesse $\mathcal{F}$. The cavities are aligned along the direction of rotation $\vec{\Omega}$. We produce linearly polarized light from a suitable source. This light is split and one part of it is sent to a suitable optical element [i.e., a quarter-wave ($\lambda/4$) plate] which converts it to left-circularly polarized light while the other part is sent to a different quarter-wave plate that converts that part to right-circularly polarized light. The two beams are sent to two different Fabry–Pérot cavities where they are stored for a time $\sim\mathcal{F}L$. After this time, the beams are sent back through the quarter-wave plates and recombined. The interfered beam is detected at the detector.
@@
In this scheme, the wavenumbers of the left- and right-circularly polarized light experience different shifts. This difference manifests as a relative phase shift between these two beams when they are re-interfered. We want the signal in this setup to scale with $\mathcal{F}L$, the total time for which the light is held in the cavity. If the mirrors of the cavity are made with conventional reflecting surfaces, the overall phase of the electromagnetic vector potential $\left(A_{\pm}\right)_{\mu}$ changes by $\pi$ upon reflection. But, the reflection does not cause a relative phase between the $x$ and $y$ components of $\left(A_{\pm}\right)_{\mu}$. It can then be shown that for the reflected wave also the wavenumbers are $k_{\pm}=\omega\mp\Omega$. Thus, the phase accumulated by the light will continually add. In other words, while left- (right-) circularly polarized light becomes right- (left-) circularly polarized upon reflection, since the direction of the light has also changed, the effect of the rotation on the wavenumber remains the same. [^2] The situation is thus similar to the accumulation of phase for light stored in a cavity in the presence of a static axion gradient (cf. [15]), as opposed to a long-period time-dependent dark-matter axion field for which the phase accumulation cancels out when employing conventional mirrors in the cavity (cf. [12]). Another way to understand the continual addition of phase regardless of the direction of travel of the light in the cavity is that the flip of left- to right-handedness (and vice versa) of the light upon reflection off the mirrors is an helicity flip that occurs because the linear momentum of the light changes sign but the angular momentum does not. Because the relative orientation of the angular momentum of the light and that of the rotation of the cavities is however unchanged after the bounce off the mirror, the phase shift to the light induced by the rotation continues to add for trips in either direction in the cavity.
@@ IV Sensitivity
The envisaged operation of this light-based gyroscope would be as a final fine-measurement stage. That is, we presume that this gyroscopic assembly would be mounted on a set of nulling gimbals designed to cancel most of the gross rotational motion of the platform being monitored. It is for this reason that we have analyzed only rotational motion oriented around the axis of the light propagation; large rotational motion around another axis would induce additional phase shifts and phase gradients across cavity mirrors that would complicate the analysis. If fine measurement along multiple axes is required, either a set of orthogonal light gyroscopes could be used, or the additional phase shifts induced by rotation about an axis different from the cavity symmetry axis would need to be understood in more detail. See also Appendix B for further discussion.
@@ Acknowledgements.
Note Added: After the appearance of the arXiv v1 preprint of this work, we became aware of [16], which discusses the effect of rotation on the polarization of light, including the frequency independence of the signal. But, [16] does not exploit the fact that the frequency independence of the signal can be used to implement this measurement strategy using RF light instead of optical light. The use of RF light considerably enhances the sensitivity of the measurement scheme since it can be stored for much longer time than optical light, enabling the setup to reach sensitivities comparable to a Sagnac configuration that is shot-noise limited. Moreover, [16] also does not exploit the frequency independence of the signal to suppress phase shifts arising from vibrations, one of the major sources of noise in conventional Sagnac gyroscopes.
Appendix A Rotations around the $\hat{z}$ axis
In this appendix we derive the results of Eqs. (2) and (3). Let us consider Maxwell’s equations in a frame rotating around the $\hat{z}$ axis $\vec{\Omega}=(0,0,\Omega)$. We have the following metric:
$g_{\mu\nu}dx^{\mu}dx^{\nu}=[-1+\Omega^{2}(x^{2}+y^{2})]dt^{2}+2\Omega dt(xdy-ydx)+dx^{2}+dy^{2}+dz^{2}\>.$ (7)
Maxwell’s equations will be $\nabla_{\mu}F^{\mu\nu}=0$, where $\nabla_{\mu}$ is a covariant derivative with respect to $g_{\mu\nu}$. Looking for plane-wave solutions propagating along $z$, we can pick Weyl gauge and take the following Ansatz for the vector potential:
$A_{\mu}=\big{[}0,A_{x}(z,t),A_{y}(z,t),A_{z}(x,y,z,t)\big{]}\>.$ (8)
As we show here, this Ansatz solves Maxwell’s equations in the rotating frame. Indeed, computing $\nabla_{\mu}F^{\mu\nu}=0$, we find the following:
$\begin{split}&\partial_{t}\partial_{z}A_{z}+\Omega\left[\partial_{z}^{2}(xA_{y}-yA_{x})+(y\partial_{x}-x\partial_{y})\partial_{z}A_{z}\right]=0\>,\\ &(\partial_{t}^{2}-\partial_{z}^{2})A_{x}+\partial_{x}\partial_{z}A_{z}\\ &\quad\quad\quad=\Omega(\partial_{t}A_{y}+y\partial_{t}\partial_{z}A_{z})+\Omega^{2}\left[xy\partial_{z}(\partial_{z}A_{y}-\partial_{y}A_{z})-y^{2}\partial_{z}(\partial_{z}A_{x}-\partial_{x}A_{z})\right]\>,\\ &(\partial_{t}^{2}-\partial_{z}^{2})A_{y}+\partial_{y}\partial_{z}A_{z}\\ &\quad\quad\quad=-\Omega(\partial_{t}A_{x}+x\partial_{t}\partial_{z}A_{z})+\Omega^{2}\left[xy\partial_{z}(\partial_{z}A_{x}-\partial_{x}A_{z})-x^{2}\partial_{z}(\partial_{z}A_{y}-\partial_{y}A_{z})\right]\>,\\ &(\partial_{t}^{2}-\partial_{x}^{2}-\partial_{y}^{2})A_{z}-\Omega\partial_{t}\left[\partial_{z}(yA_{x}-xA_{y})+2(x\partial_{y}-y\partial_{x})A_{z}\right]\\ &\quad\quad\quad=\Omega^{2}\left[x(\partial_{x}A_{z}-\partial_{z}A_{x})+y(\partial_{y}A_{z}-\partial_{z}A_{y})-(x^{2}\partial_{y}^{2}+y^{2}\partial_{x}^{2}-2xy\partial_{x}\partial_{y})A_{z}\right]\>.\end{split}$ (9)
These equations can be solved with the following choice:
$A_{\mu}=e^{i(\omega t-k_{\sigma}z)}\big{[}0,1,i\sigma,\alpha(x+i\sigma y)\big{]}\qquad[\sigma=\pm 1]\>.$ (10)
Indeed, plugging this choice in the first Maxwell equation, this becomes (using $\sigma^{2}=1$):
$\begin{split}(x+i\sigma y)[\alpha\omega-i\sigma\Omega(k_{\sigma}-i\alpha)]k_{\sigma}=0\>,\end{split}$ (11)
which is satisfied for $k=\omega-\sigma\Omega$ and $\alpha=i\sigma\Omega$, confirming Eq. (3). Similarly, the other three (spatial) Maxwell equations become (again using $\sigma^{2}=1$):
$\begin{split}\omega^{2}-k_{\sigma}^{2}+i\alpha k_{\sigma}-\omega\sigma\Omega&=-(x+i\sigma y)\left[\alpha\omega-(ik_{\sigma}+\alpha)\sigma\Omega\right]yk_{\sigma}\Omega\;,\\ -i\sigma(\omega^{2}-k_{\sigma}^{2}+i\alpha k_{\sigma}-\omega\sigma\Omega)&=-(x+i\sigma y)\left[\alpha\omega-(ik_{\sigma}+\alpha)\sigma\Omega\right]xk_{\sigma}\Omega\;,\\ (x+i\sigma y)[\alpha(\omega-\sigma\Omega)^{2}&-i\sigma\Omega k_{\sigma}(\omega-\sigma\Omega)]=0\>.\end{split}$ (12)
One can easily verify that these equations are solved as well by the choice $k=\omega-\sigma\Omega$, $\alpha=i\sigma\Omega$.
Appendix B Generic rotations
Our purpose in this appendix is to demonstrate that the phase shift of the transverse vector potentials considered in the main text (and derived in detail in appendix A) still holds even when there are additional small rotations about axes orthogonal to the cavity symmetry axis.
@@ Appendix B Generic rotations
we can perturb about the solution at (2), which is exact when $\Omega_{x}=\Omega_{y}=0$. Suppose that we treat $\Omega_{x},\Omega_{y}$ as formal small parameters by sending $\Omega_{x,y}\rightarrow\epsilon\Omega_{x,y}$, and that we write each component of the vector potential as a power series in $\epsilon$. Substituting into Maxwell’s equations (MEs), expanding in powers of $\epsilon$, and satisfying the resulting equations that occur at each independent power of $\epsilon$ separately, allows us to derive a systematic perturbative expansion in powers of the formal small parameter $\epsilon$. Moreover, we will also consider the solution only close to the symmetry axis of the cavity, such that we can take $x,y\rightarrow 0$ after substituting into MEs. [^3] Of course, because MEs are differential equations, it is important for the purposes of checking if this is a good solution to keep terms $\sim x^{n},y^{m}$ until after substitution in the MEs. That is, to only set $x,y\rightarrow 0$ after such substitution.
@@
We therefore conclude that, as long as the mild conditions derived in this appendix hold, the approach we advanced in the main text will continue to apply even when $\Omega_{x,y}\neq 0$, so long as these rotation rates are sufficiently small.
Appendix C Inertial frame and rotation of the apparatus
@@
$A_{\mu}=e^{i[(\omega-\sigma\Omega)t-kz]}(0,1,i\sigma,0)\quad,\quad k=\omega-\sigma\Omega\;.$ (16)
@@
This makes clear that reading the signal from a rotating laboratory does not remove the polarization-dependent phase shift. Indeed, while the time-dependent part of the phase shift between different polarizations is absorbed by the basis, the space-dependent relative phase accumulated over the interval $L$ does not cancel. This is the same phase shift one sees in the non-inertial frame.
@@
$2A_{\mu}=e^{i[(\omega-\Omega)t-k_{+}z]}(0,1,i,0)+e^{i[(\omega+\Omega)t-k_{-}z]}(0,1,-i,0)\quad,\quad k_{\pm}=\omega\mp\Omega\;.$ (18)